OpenStudy (anonymous):

determine wether mn(line) and rs(line) are perpindicular parrallel or neither m(-1,3) N(0,5) r(2,1) s(6,-1)

OpenStudy (anonymous):

they are neither parallel nor perpindicular

OpenStudy (sriram):

agree wid rajen slopes are 1/2 & -1/2

OpenStudy (anonymous):

there slopes of mn is 2 rs is -1/2

OpenStudy (anonymous):

phi i give u a link some1 answer

OpenStudy (phi):

two lines are parallel if they have identical slopes (though be careful, they could be the same line). two lines are perpendicular if m1= -1/m2, where m1, m2 are the slopes of line 1 and line 2, respectively. The slope of line mn is (change in y)/(change in x) using m(-1,3) N(0,5), we get \[\frac{5 - 3}{0- -1}=2 \] the slope of line rs, using r(2,1) s(6,-1), is \[\frac{-1 -1}{6-2}= \frac{-2}{4}= -\frac{1}{2} \] Because 2= -1/2, the lines are perpendicular.